OpenStudy (anonymous):

Before making invitations, Katrina counted the supplies she bought for her party. She counted 36 water balloons and 81 beads to make bracelets. She then needed to decide how many invitations to make. What is the largest number of friends she can invite so each person gets the same amount of water balloons and each friend gets the same number of beads?

OpenStudy (anonymous):

o and i doing a big test to

OpenStudy (anonymous):

waht !!!!!!!!!

OpenStudy (liv1234):

Alright, so to start off, can you think of a number that goes into 36 and 81?

OpenStudy (liv1234):

For example, 5 goes into 20 and 10.

OpenStudy (liv1234):

I have the answer, but I wanna make sure you understand why it is the answer. :p

OpenStudy (anonymous):

9 is the highest integer which will divide into 36 and 81 a whole number of times (assuming this is what you actually want to know). 36 = 9 * 4 81 = 9 * 9 Of course, we could also answer this specific question with "1" or "3" also.

OpenStudy (liv1234):

Correct! I was hoping you said 9.

OpenStudy (anonymous):

lloll

OpenStudy (anonymous):

so the answer is 9

OpenStudy (liv1234):

We have to make sure that the amount of friends she has are able to all get an equal amount of water balloons and beads, though.

OpenStudy (anonymous):

here are the choices

OpenStudy (liv1234):

Okay because seeing the options will make it easier to solve.

OpenStudy (anonymous):

A) 6 B) 7 C) 9 D) 10

OpenStudy (liv1234):

No, it's most definitely 9. I apologize, I was getting confused.

OpenStudy (anonymous):

o ok